# InSyn: Modeling Complex Interactions for Pedestrian Trajectory Prediction

### v1

@@ Abstract
Accurate pedestrian trajectory prediction is crucial for intelligent applications, yet it remains highly challenging due to the complexity of interactions among pedestrians. Previous methods have primarily relied on relative positions to model pedestrian interactions; however, they tend to overlook specific interaction patterns such as paired walking or conflicting behaviors, limiting the prediction accuracy in crowded scenarios. To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors. Additionally, we introduce a training strategy termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction. Experiments on the ETH and UCY datasets demonstrate that our model outperforms recent baselines significantly, especially in high-density scenarios. Furthermore, the SSOS strategy proves effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
[…]
7556
1 Introduction
@@
Pedestrian trajectory prediction plays a crucial role in various fields such as autonomous driving [21], robotic navigation [4], and smart cities [5]. Accurately forecasting a pedestrian’s future trajectory not only helps intelligent systems better understand human behavior but also significantly enhances the safety and efficiency of human-machine interaction. However, beyond an individual’s own motion, interactions among pedestrians are often highly complex, making trajectory prediction a challenging task. In a given scene, pedestrians may respond to nearby individuals in different ways. For instance, the trajectories of two pedestrians may exhibit conflict. Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group. Notably, even in walk-in-sync scenarios, subtle interactions may still occur. Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
@@
Figure 1: Comparison of Interaction Modeling: Previous Methods vs. Our Approach. F represents the interaction affect between pedestrians. In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors. This simplistic assumption may lead to overestimating the influence of fellow pedestrians. Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians (e.g., In Sync or Conflict), providing a more nuanced understanding of the neighboring interaction.
@@
Recently, extensive studies have investigated pedestrian interactions [1, 32, 25, 43], using approaches like social pooling layers [1, 7], social force mechanism [43], and graph neural networks (GNNs) [41, 19]. However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, without considering the specific interaction relationships between them. This limitation restricts the capacity of previous models to capture the dynamics of real-world interactions, such as walking in sync or avoiding collisions, which are crucial for accurately predicting trajectory in crowded scenario. Prior approaches treat all neighboring pedestrians uniformly, failing to differentiate patterns such as In Sync and Conflict (see Figure 1). This oversimplification may leads to overfitting on sparse datasets and underperforming in complex social scenarios. For instance, in situations where pedestrians walk in sync (e.g., friends walking side-by-side), failing to account for this coordinated behavior may lead to overlapping trajectories during prediction. Similarly, in conflict scenarios (e.g., two pedestrians approaching each other head-on or navigating narrow passages), the lack of explicit modeling for competitive or avoidance behaviors often results in unrealistic trajectories that either overestimate collision risk or underestimate the necessity for trajectory alteration. These limitations become particularly pronounced in crowded or dynamic environments.
To overcome this issue, we need a method that explicitly identifies and captures these interaction patterns, enabling a deeper understanding of the diverse pedestrian interactions, as shown in Figure 1.
@@
Although pedestrian trajectories are highly stochastic, they usually have a clear goal [28]. By focusing on the target goal, the performance of the model can be significantly enhanced. Experimental results from Giuliari et al. [9] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goal), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
@@
This paper proposes the InSyn, a Transformer-based model. The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism. (2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step. (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goals sampling.
@@
We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [22] and UCY [16]. The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions. We further conduct ablation studies to show the superiority of interaction modeling and SSOS strategy.
@@
  • We propose a pedestrian interaction modeling approach that explicitly identifies and leverages specific interaction patterns, achieving a significant improvement in average ADE compared to the previous baselines and offering more comprehensive insights into socially-aware motion patterns.
@@
2 Related Work
2.1 Pedestrian Trajectory Prediction
Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches. Traditional methods, such as Markov decision processes [30] and Gaussian distributions [35], are efficient in handling simple scenarios. However, they often struggle with complex data distributions and exhibit limited generalization capabilities. In contrast, deep learning approaches have shown significant improvements in modeling complex patterns [13, 8, 6, 36]. Long Short-Term Memory (LSTM) network [11], a classic time-series model, is applied in this field. Social-LSTM [1] introduced a pooling mechanism to incorporate interactions into trajectory prediction. Yang et al. [40] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [44] combined neighbor intentions and message passing to enhance prediction accuracy. These work leverage the strong temporal feature extraction capabilities of LSTM.
In recent years, Transformer have gained popularity in trajectory prediction due to their ability to capture long-range dependencies. Liu et al. [20] combined CNN and Transformer, using attention mechanism to process interaction information. MRGTraj [23] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure. CITraNet [3] significantly improved prediction accuracy by introducing an innovative Transformer-based Gumbel distribution network.
Building on these advancements, our research combines the strengths of LSTM and Transformer. By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interactions features, thereby improving prediction performance.
2.2 Pedestrian Interaction Modeling
@@
Current mainstream methods for interaction modeling include social pooling layers [1, 37], GNNs [19, 41], and attention mechanisms [18, 38]. Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [37]. Yang et al. [38] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy. Transformer have been utilized in pedestrian interaction modeling [34, 39, 12, 2, 27]. For example, Yuan et al. [42] proposed AgentFormer that learns spatiotemporal interaction embeddings from sequential trajectory features. Similarly, Yang et al. [39] leveraged GNNs and Transformer to model spatial and temporal dependencies. To address future interaction modeling, Amirloo et al. [2] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
However, most of these methods fail to consider the specific interaction relationships among pedestrians, such as walking In Sync or engaging in Conflict. To address this limitation, we propose a model that explicitly captures specific types of interaction features.
3 Problem Formulation
@@
where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions respectively. These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2 .
Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time $t-1$, the interaction state is classified as In Sync; otherwise, it is marked as Conflict. In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible. However, if a new pedestrian suddenly enters in a region, referred as Conflict, it is more likely to impact the agent’s trajectory. The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region. A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence. The goal of this study is to train a generative model
@@
where $\theta$ represents the model parameters. The model aims to predict the distribution of pedestrian’s future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
Figure 3: Overview of the InSyn framework for trajectory prediction. Our model consists of three key modules: (1) Interaction Encoder, (2) Trajectory Generator, and (3) Seq-CVAE. The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the obserbed time $0:\tau$.
4 Methodology
4.1 Framework Design
In this section, we propose InSyn, a Transformer-based model designed to identify and model spefific social interaction. The model consists of three components: (1) Interaction Encoder: Extracts interaction features and enables goal-driven prediction. (2) Trajectory Generator: Employing SSOS strategy to mitigate initial prediction step divergence. (3) Seq-CVAE Goal Sampler: A generator for sequential data that utilizes high-dimensional latent variables for goal sampling. The framework of InSyn is illustrated in Figure 3, and the Seq-CVAE is depicted in Figure 5.
@@ 4.2 Interaction Encoder
4.2.1 Goal-Driven
Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  4.3) is combined with the position information through concatenation. Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions. This encoding is based on the sequence positions of the inputs. Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
4.2.2 Interaction
The complex interactions among pedestrians are extracted through a neighbor encoder module follwed by an LSTM. The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder. The design of the neighbor encoder includes three components: the embedding layer, the Distance Gate, and the Social Pooling Layer. The embedding layer maps three interaction states (No Neighbor, In Sync, Conflict) into corresponding representations. These representations are then combined with the processed distance $d$ through element-wise multiplication. The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer. The Distance Gate is similar to the gate mechanism in LSTM [11]. It maps $d$ to a weight in the range [0,1]. This weight reflects distance-based intensity, enhancing the spatial characteristics of pedestrian interactions.
@@ 4.2.2 Interaction
4.3 SSOS: Seq-Start of Seq
The Trajectory Generator incorporates the SSOS Strategy into the Transformer Decoder [29] to alleviate initial prediction divergence. The Transformer’s decoder relies on the Start of Sequence (SOS) as the initial input when generating the first output. In Natural Language Processing (NLP), the Begin-of-Sentence token $bos\right>$ is commonly used as the SOS, which aligns with the semantic structure of text. However, in sequential trajectory prediction, the selection of SOS should consider the spatial information; otherwise, it may introduce noise. Giuliari et al. [9] set the SOS to $(0,0)$ for trajectories prediction. While $(0,0)$ represents a fixed position in spatial coordinates, using it as the SOS may introduce noise and mislead the model due to its lack of alignment with the observed trajectory. Therefore, it is reasonable to consider using the position at the last observed time step ,$\widehat{pos}_{\tau}$, as the SOS token of decoder.
Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$. That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$. This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with SOS and the cross-attention with the encoder’s output. Insufficient information or excessive noise in the SOS may lead alignment bias, thus reducing prediction accuracy.
Figure 4: Comparison of SSOS and SOS.
To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS. We refer to this approach as SSOS (Figure 4) and use it during training. In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent. The loss function using SSOS is presented below:

### v2

# Trustworthy Pedestrian Trajectory Prediction via Pattern-Aware Interaction Modeling

@@ Abstract
Accurate and reliable pedestrian trajectory prediction is critical for the safety and robustness of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians. Previous methods often adopt black-box modeling of pedestrian interactions, treating all neighbors uniformly. Despite their strong performance, such opaque modeling limits the reliability of predictions in safety-critical real-world deployments. To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors. Additionally, we introduce a training strategy, termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction. Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides stronger interpretability, achieving a favorable trade-off between reliability and accuracy. Furthermore, the SSOS strategy proves to be effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
Index Terms:
Human Trajectory Prediction, Interpretable Modeling, Attention Mechanism
I Introduction
@@
Pedestrian trajectory prediction is essential for safety-critical applications such as autonomous driving [1] and robotic navigation [2]. Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures safe and trustworthy predictions. However, the presence of complex pedestrian interactions poses significant challenges to the task. In a given scene, pedestrians may respond to nearby individuals in different ways. For instance, the trajectories of two pedestrians may exhibit conflict. Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group. In particular, even in walk-in-sync scenarios, subtle interactions may still occur. Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
@@
Figure 1: Comparison of Interaction Modeling: Previous Methods vs. Our Approach. F represents the interaction effect between pedestrians. In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors. This simplistic assumption may lead to overestimating the influence of fellow pedestrians. Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians (e.g., In Sync or Conflict), providing a more nuanced understanding of the neighboring interaction.
@@
Recently, extensive studies have investigated pedestrian interactions [3, 4, 5, 6], using approaches like social pooling layers [3, 7], social force mechanism [6], and graph neural networks (GNNs) [8, 9]. However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1). This oversimplification may lead to overfitting and untrustworthy predictions. For example, ignoring synchronized walking (e.g., friends walking side-by-side) can lead to overlapping predicted paths. In conflict cases, such as head-on encounters, the absence of explicit modeling may cause unrealistic predictions that either overestimate collisions or miss necessary evasive actions. This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in safety-critical scenarios such as machine navigation and intelligent transportation systems. These limitations become particularly pronounced in crowded or dynamic environments.
To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing concrete interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and interpretable predictions (see Figure 1).
@@
Although pedestrian trajectories are highly stochastic, they usually have a clear goal [10]. By focusing on the target goal, the performance of the model can be significantly enhanced. Experimental results from Giuliari et al. [11] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goals), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
@@
This paper proposes the InSyn, a Transformer-based model. The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism. (2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step. (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goal sampling.
@@
We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [12] and UCY [13]. The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions. We further conduct ablation studies to show the superiority of interaction modeling and the SSOS strategy. Additionally, the case study highlights the interpretability of our approach by illustrating how specific interaction patterns influence prediction outcomes.
@@
  • We propose a pedestrian interaction modeling approach that explicitly identifies and leverages specific interaction patterns, achieving a significant improvement in average ADE compared to the previous black-box baselines and contributing to more trustworthy socially-aware trajectory prediction.
@@
II Related Work
II-A Pedestrian Trajectory Prediction
Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches. Traditional methods, such as Markov decision processes [14] and Gaussian distributions [15], are efficient in handling simple scenarios. However, they often struggle with complex data distributions and exhibit limited generalization capabilities. In contrast, deep learning approaches have shown significant improvements in modeling complex patterns [16, 17, 18, 19]. Long Short-Term Memory (LSTM) network [20], a classic time-series model, is applied in this field. Social-LSTM [3] introduced a pooling mechanism to incorporate interactions into trajectory prediction. Yang et al. [21] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [22] combined neighbor intentions and message passing to enhance prediction accuracy. These works leverage the strong temporal feature extraction capabilities of LSTM.
In recent years, Transformers have gained popularity in trajectory prediction due to their ability to capture long-range dependencies. Liu et al. [23] combined CNN and Transformer, using the attention mechanism to process interaction information. MRGTraj [24] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure. CITraNet [25] significantly improved prediction accuracy by introducing an innovative Transformer-based Gumbel distribution network.
Building on these advancements, our research combines the strengths of LSTM and Transformer. By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interaction features, thereby improving prediction performance.
II-B Pedestrian Interaction Modeling
@@
Current mainstream methods for interaction modeling include social pooling layers [3, 26], GNNs [9, 8], and attention mechanisms [27, 28]. Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [26]. Yang et al. [28] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy. Transformers have been utilized in pedestrian interaction modeling [29, 30, 31, 32, 33]. For example, Yuan et al. [34] proposed AgentFormer that learns spatiotemporal interaction embeddings from sequential trajectory features. Similarly, Yang et al. [30] leveraged GNNs and Transformers to model spatial and temporal dependencies. To address future interaction modeling, Amirloo et al. [32] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
However, most of these methods rely on black-box representations of social influence. This limits the model’s ability to produce reliable and interpretable predictions, especially in complex or safety-critical scenarios. In contrast, our approach explicitly captures structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
III Problem Formulation
@@
where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions, respectively. These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2. The construction of interaction patterns is highly transparent and trustworthy, as it relies on interpretable spatial-temporal neighbor changes.
Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time t1t-1, the interaction state is classified as In Sync; otherwise, it is marked as Conflict. In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible. However, if a new pedestrian suddenly enters a region, referred to as Conflict, it is more likely to impact the agent’s trajectory. The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region. A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence. The goal of this study is to train a generative model
@@
where $\theta$ represents the model parameters. The model aims to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
Figure 3: Overview of the InSyn framework for trajectory prediction. Our model consists of three key modules: (1) Interaction Encoder, (2) Trajectory Generator, and (3) Seq-CVAE. The input observed walking state includes the agent’s trajectory positions $pos_{0:\tau}$ and its interaction information $N_{0:\tau}$ within the observed time $0:\tau$.
IV Methodology
In this section, we propose InSyn, a Transformer-based model tailored for the interpretable modeling of specific social interactions. The model consists of three components: (1) Interaction Encoder: Extracts interaction features and enables goal-driven prediction. (2) Trajectory Generator: Employing the SSOS strategy to mitigate initial prediction step divergence. (3) Seq-CVAE Goal Sampler: A generator for sequential data that utilizes high-dimensional latent variables for goal sampling. The framework of InSyn is illustrated in Figure 3, and the Seq-CVAE is depicted in Figure 4.
@@ IV-A Interaction Encoder
IV-A1 Goal-Driven
Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  IV-B) is combined with the position information through concatenation. Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions. This encoding is based on the sequence positions of the inputs. Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
IV-A2 Interaction
To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware and distance-based interaction effects.
The interaction patterns between pedestrians are extracted through a neighbor encoder module followed by an LSTM. The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder. which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer. The embedding layer maps interaction states (No Neighbor, In Sync, Conflict) to learnable representations. These representations are then combined with the processed distance $d$ through element-wise multiplication. The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer. Inspired by the gating mechanism in LSTM [20], the Distance Gate maps $d$ to a weight in [0,1], reflecting distance-based interaction intensity.
@@ IV-A2 Interaction
IV-B SSOS: Seq-Start of Seq
[…]
Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$. That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$. This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with the SOS and the cross-attention with the encoder’s output. Insufficient information or excessive noise in the SOS may lead to alignment bias, thus reducing prediction accuracy.
To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS. We refer to this approach as SSOS and use it during training. In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent. The loss function using SSOS is presented below: